a) In how many ways can a committee of 5 be chosen from 9 people?



(b) How many committees of 5 or more can be chosen from 9 people?



(c) In how many ways can a committee of 5 teachers and 4 students be chosen



from 9 teachers and 15 students?



(d) In how many ways can the committee in (C) be formed if teacher A refuses



to serve if student B is on the committee?

1
Expert's answer
2021-12-28T13:25:55-0500

a) If the order matters (we distinguish people of a committee)


"\dfrac{9!}{(9-5)!}=9(8)(7)(6)(5)=15120"

If the order does not matter

"\dbinom{9}{5}=\dfrac{9!}{5!(9-5)!}=\dfrac{9(8)(7)(6)(5)}{120}=126"


(b) If the order matters


"\dfrac{9!}{(9-5)!}+\dfrac{9!}{(9-6)!}+\dfrac{9!}{(9-7)!}"

"+\dfrac{9!}{(9-8)!}+\dfrac{9!}{(9-9)!}"

"=15120+60480+181440+362880+362880"

"=982800"

If the order does not matter

"\dbinom{9}{5}+\dbinom{9}{6}+\dbinom{9}{7}+\dbinom{9}{8}+\dbinom{9}{9}""=126+84+36+9+1=256"

(c)


"\dbinom{9}{5}\dbinom{15}{4}=\dfrac{9!}{5!(9-5)!}\cdot \dfrac{15!}{4!(15-4)!}"

"=126\cdot1365=171990"

(d)

In how many ways can a committee of 5 teachers and 4 students be chosen

from 9 teachers and 15 students if  teacher A and student B are on the committee

teacher A: "\dbinom{8}{4}"

student B: "\dbinom{14}{3}"

 teacher A and student B "\dbinom{8}{4}\dbinom{14}{3}"

Then the number of ways can the committee in (C) be formed if teacher A refuses to serve if student B is on the committee is



"\dbinom{9}{5}\dbinom{15}{4}-\dbinom{8}{4}\dbinom{14}{3}"

"=171990-70(364)=146510"